OpenStudy (anonymous):

What is the pattern for these numbers? 11 33 64 77 88 101 (Calculus 2)

myininaya (myininaya):

I'm trying to think about your pattern here. lol

myininaya (myininaya):

I think there is a way to do it on the calculator, but I never learned how to.

myininaya (myininaya):

Okay, help me think. If x=11, then 3x=33, 2(3x)-2=64, 3(3x)-2(10)-2=77, 4(3x)-2(10)-2(10)-2-2=88, and 5(3x)-2(10)-2(10)-2(10)-2-2=101

OpenStudy (anonymous):

Ok, I'm confused now... I get that algebra to all this, but how did you get that? What is the pattern?

myininaya (myininaya):

i still don't know the pattern. What have you learned about patterns? I was trying to think really really hard, and maybe over thinking.

OpenStudy (anonymous):

Yes i was thinking the same thing... I will just ask for an explanation tomorrow.

myininaya (myininaya):

I would like to know

myininaya (myininaya):

If you remember, tell me

myininaya (myininaya):

It drives me crazy when I can't figure something out

OpenStudy (anonymous):

Ditto!